Suppose you are a stock analyst at Goldman Sachs. You forecast that Paddy plc will pay dividends of $2.50, $2.75, and $2.25 at the end of each of the next 3 years. It is expected that thereafter the firm will sustain a long-run rate of growth of 4.5%. The appropriate cost of equity for Paddy is 10%. Based on your forecasts, what is the “fair” price for Paddy?
Answer #1

Let Dividend in Year n be denoted by Dn

Given, D1 = $2.50

D2 = $2.75

D3 = $2.25

Growth rate for subsequent years = g = 4.5%

Hence, D4 = D3(1+g) = 2.25(1+0.045) = $2.35

Cost of Equity = r = 0%

Using Gordons Growth model,

Price of Stock in Year 3 = P3 = D3/(r - g) = 2.35/(0.10 - 0.045) = $42.73

Hence, Price of stock now = P0 = D1/(1+r) + D2/(1+r)2 + D3/(1+r)3 + P3/(1+r)3

= 2.50/(1+0.10) + 2.75/(1+0.10)2 + 2.25/(1+0.10)3 + 42.73/(1+0.10)3

= $38.34